# Efficient Monte-Carlo sampling of metastable systems using non-local collective variable updates

### v2

@@ III.3 Dimer in a solvent
We now apply the algorithm on a model with a non-linear collective variablestraub_molecular_1988; dellago_calculation_1999; lelievre-rousset-stoltz-book-10. It consists of $N$ particles of equal masses $M$ in a two-dimensional box of side length $L$ with periodic boundary conditions. Two of the particles form a dimer; the remaining particles are designated as solvent particles. The solvent particles interact between them and with the dimer particles via the Weeks–Chandler–Andersen (WCA) pair potential, which is a truncated Lennard-Jones potential:
@@
The minima of $V_{\mathrm{D}}$ are at $r=r_{0}$ and $r=r_{0}+2w$, giving rise to a compact state and stretched state with an energy barrier of height $h$. Two configurations for the compact and the stretched state are shown in Fig.˜7.
@@
We ran numerical experiments for parameters $h=w=2$, $r^{*}=\epsilon=1$ and $M=1$ with $N=100$ particles and a box of length $L=15$, in line with previous workslelievre-rousset-stoltz-book-10; lelievre_langevin_2012. For the CV sampler we use a Gaussian mixture model with modes of equal weight at $z=0$ and $z=1$ and width $\sigma_{\mathrm{prop}}=0.2$. We reject any unphysical proposals with $z<z_{\mathrm{min}}=-\frac{r_{0}}{2w}$ and $z>z_{\mathrm{max}}=(L/\sqrt{2}-r_{0})/(2w)$ since the distance $|q_{1}-q_{2}|$ is always positive and the periodic boundary conditions (torus coordinates) impose a maximum possible distance.
We compare the performance of the algorithm for different parameter values using a mode jump cost that counts the number of calls to the force required to go from a state with $z<0.1$ to a state with $z>0.9$ or vice versa, in line with the ‘mean residence duration’ used in Ref. lelievre-rousset-stoltz-19.
Once again, we observe a large advantage in using the deterministic algorithm with $\alpha_{1}=0$, whose performance for different values of $\alpha_{2}$ and $\widetilde{v}$ are shown in Fig.˜8, while we see almost no mode switches for the overdamped algorithm within the same computational budget (see Section˜H.4).
In Fig.˜9, we show the free energy profile obtained by running the algorithm for the optimized parameters until convergence together with the free energy in the absence of solvent particles (see Section˜F.3 for its derivation). As one can see, the presence of the solvent particles increases the likelihood of the compact state (lowering its free energy in comparison with the stretched state). The result agrees almost perfectly with a result obtained from thermodynamic integration (TI) (see Ref. lelievre_langevin_2012 for a theoretical introduction). In thermodynamic integration, the system is sampled under the constraint $\xi(q)=z_{i}$ for a sequential series of CV-values $z_{i}$ on a grid to estimate the mean force. This demonstrates that the inaccuracy of the proposal potential has been corrected by the algorithm.
Note that this model was also considered by the authors of Ref. nilmeier2011nonequilibrium, Contrary to our approach, however, they chose the full dimer configuration as the CV, $\xi(q)=(q_{1},q_{2})^{\top}$, and used a proposal kernel that samples a dimer configuration $\widetilde{Z}_{n+1}$ conditioned on the current configuration $Z_{n}$ with a new distance $|q_{1}-q_{2}|$ but the same overall rotation angle.
@@
Figure 9: Free energy profile for the dimer model obtained from running the CV guided algorithm with the optimal parameters determined in Fig.˜8 with 20 chains and 50,000 iterations, shown alongside results from thermodynamic integration on a grid of 50 values and $10^{6}$ steps each. For comparison, we also show the free energy profile of the dimer without solvent particles, see Section˜F.3, and the energy of the proposal distribution. An energy shift is applied to all of these curves such that they coincide at $z=1$. Error bars are estimated from the standard deviation over parallel MCMC chains of counts of the histogram bins.
@@ III.4 Polymer in a solvent
As a final example, we consider a toy molecular system that consists of a 9-bead polymer embedded in a three-dimensional box with 200 solvent particles, similar to a model introduced in Ref. tamagnoneCoarseGrainedMolecularDynamics2024. The polymer beads and the solvent particles all interact via a Lennard-Jones (LJ) potential. Within the polymer, the nearest and second-nearest neighbor distances (and hence the binding angle) of the polymer beads are softly constrained with a harmonic potential. In addition, a bi-modal potential is imposed on the end-to-end distance of the polymer, giving rise to two stable confirmations in an open and a closed state, see Fig.˜10. Denoting by $\{q_{i}\}_{i=1,\dots,9}$ the polymer coordinates and by $\{q_{i}\}_{i=10,\dots,N}$ the solvent coordinates, the total energy is given by
@@
Figure 11: Simulation of the polymer with a one-dimensional CV (end-to-end distance). Left panel: Rough grid optimization of parameters for the deterministic case $\alpha_{1}=0$, each simulation run with 20 parallel chains and 2,000 iterations. Right panel: Marginals for a run of 20 chains and 100,000 iterations for the optimized parameters $1/\widetilde{v}=180$, $\alpha_{2}=2^{-4}$ (overall 8% acceptance), shown alongside the CV proposal distribution and reference samples obtained from a biased simulation that has been reweighted.
@@
An obvious choice of CV that resolves the bimodality is the end-to-end distance. Since it is one-dimensional, it is easy to construct a proposal distribution in the form of a Gaussian mixture. For simplicity, we extract the position of the modes and their width from short Langevin simulations on the full system (see Appendix˜G). A rough grid search to optimize the parameters of the algorithm is shown in Fig.˜11, with the best achieved inverse mode jump cost at around $9\times 10^{-5}$. In contrast, the overdamped algorithm with $\alpha_{1}=1$ does not show any transitions within the same parameter range and computational budget, giving an upper limit of the performance of at most $6\times 10^{-7}$; it is thus at least two orders of magnitude worse in performance.
Another possible choice of collective variable in this case is to use the positions of the beads of the entire polymer configuration. Here, a good proposal distribution requires a careful parameterization, for which we use a normalizing flow (see Appendix˜G). Since our primary interest is in the performance of the algorithm, we trained our proposal sampler on already existing reference data. In practice, one could use adaptive schemes in which the CV-sampler is progressively trained while running the MCMC samplergabrieAdaptiveMonteCarlo2022; tamagnoneCoarseGrainedMolecularDynamics2024. Note that for this choice of CV, the steered dynamics in Step 2 of the algorithm effectively only acts on the solvent particles, since the schedule already imposes the evolution of the entire polymer. Running the algorithm, we again observe that the deterministic dynamics ($\alpha_{1}=0$) gives good results with frequent mode switches while the overdamped version ($\alpha_{1}=1$) does not work at all. Strikingly, for the optimized parameters, we achieve an overall acceptance rate of 99% with this algorithm. Since the proposals are drawn independently, this corresponds almost to the optimum performance with maximum decorrelation within the polymer degrees of freedom. A rough grid search for the deterministic algorithm is shown in Fig.˜12 together with the marginal distribution of the end-to-end distance. Comparing the optimal mode jump cost obtained with the two choices of CV, we see an advantage of around one order of magnitude in using the positions of all polymer beads over the end-to-end distance between first and last bead as a CV. This only accounts for the calls to the force, however, and not for the effort in learning the proposal sampler in CV space.
@@ IV Conclusion and outlook
In this work, we present an MCMC algorithm with non-local proposals in a collective variable space. This CV space may be high-dimensional, provided a good proposal sampler can still be learned. We generalize related algorithms by considering the underdamped Langevin dynamics and a non-linear collective variable and we prove reversibility of the resulting scheme. We provide a detailed description of the algorithm so that it is readily implementable and we illustrate its performance on a number of examples, ranging from a toy low-dimensional example to molecular systems. We demonstrate a significant performance advantage in using Hamiltonian over overdamped steered Langevin dynamics.
@@ Appendix A Numerical implementation of the evolution of momenta
For practical purposes, one can reformulate (II.17) (and (II.19) in the same fashion) in the following equivalent form that is more suitable to numerical implementation, see Sec. 5.4.2 in Ref. lelievre_langevin_2012 for further explanations:
@@ Appendix B Proof of the reversibility of the algorithm
The objective of this section is to prove the reversibility of the algorithm introduced in Section II.2 (namely Theorem˜II.4). This is done in three steps. In Section B.1, we introduce the necessary material, in particular concerning measures on submanifolds associated with the collective variable $\xi$. Then,in Section B.2, we present fundamental properties of the steered dynamics (II.17)-(II.19), and in particular the Jazynski-Crooks equality. This equality is the cornerstone of the proof of reversibility which is finally provided in Section B.3.
@@ B.1 Introducing notation
This subsection largely uses tools and results from Ref. lelievre_langevin_2012. Let us introduce a number of quantities which we will need later on. A central quantity of interest is the free energy $F:\mathbb{R}^{\ell}\rightarrow\mathbb{R}$ associated with the target measure $\nu$ and the collective variable $\xi$, defined as (see (II.2) for the definition of $\Sigma(z)$)
$F(z)=-\frac{1}{\beta}\log\int_{\Sigma(z)}Z_{q}^{-1}e^{-\beta V(q)}\delta_{\xi(q)-z}(\mathrm{d}q).$ (B.1)
@@
The proof relies on Ref. lelievre_langevin_2012. From Sec. 2.3 therein, we restate the following two identities: By the co-area formula, one can show that
@@
where we denote by $\sigma^{M}_{\Sigma(z)}(\mathrm{d}q)$ the surface measure on $\Sigma(z)$ induced by the scalar product $\langle q,\tilde{q}\rangle_{M}=q^{\top}M\tilde{q}$ on $\mathbb{R}^{d}$, and, for a given $q\in\Sigma(z)$, by $\sigma^{M^{-1}}_{\Sigma_{v_{\xi}(q,\cdot)}(v_{z})}(\mathrm{d}p)$ the surface measure on the affine space $\Sigma_{v_{\xi}(q,\cdot)}(v_{z})$, induced by the scalar product $\langle p,\tilde{p}\rangle_{M^{-1}}=p^{\top}M^{-1}\tilde{p}$ on $\mathbb{R}^{d}$. For more precise definitions of these measures, we refer to Ref. lelievre-rousset-stoltz-book-10, Sections 3.2.1 and 3.3.2 and the references therein.
@@
Equation (B.5) follows by taking $v_{z}=0$. For the last but one equality in the series of equalities above, we use a computation from Sec. 5.1 in Ref. lelievre_langevin_2012, namely, for fixed $q$,
@@ B.2 Properties of the steered dynamics
In this section, we state a number of properties of the RATTLE scheme introduced in (II.17)-(II.19). Since we follow dynamics for the energy modified by the Fixman term, in principle we would need to put a tilde on all associated measures and quantities, but we omit it here for the sake of readability and just consider the dynamics for a general potential $V$.
@@ B.2.1 Forward steered dynamics
since the Gram tensor $G_{M}(q^{k+1})=\nabla\xi(q^{k+1})^{T}M^{-1}\nabla\xi(q^{k+1})$ is invertible by assumption. In contrast, the constraint $(C_{q})$ leads to a non-linear equation for $\lambda^{k+1/2}$, which is solved using Newton’s method. In practice, we never encountered non-convergence of Newton’s method in our numerical experiments. See however Ref. lelievre-rousset-stoltz-19 and Sec. 3.2.1 of Ref. lelievre_langevin_2012) for techniques to deal with this difficulty, using so-called reversibility checks.
@@
Notice that the velocity schedule $(v_{z}^{(Z_{n},\widetilde{Z}_{n+1})}(t_{k}))_{0\leq n\leq K_{T}}$ can in principle be chosen independently of the position schedule $(z^{(Z_{n},\widetilde{Z}_{n+1})}(t_{k}))_{0\leq k\leq K_{T}}$. It is only in the continuous-time limit that the schedule $z(t)$ directly imposes the velocity schedule as $v_{z}(t)=\dot{z}(t)$. Still, certain choices can be numerically beneficial, as explained in Ref. lelievre_langevin_2012, Sec. 5.4.1. In (II.18), the unconstrained part of the update is given by
@@
which is followed by a projection with $q^{k+1}=\widetilde{q}^{k+1}+\Delta tM^{-1}\nabla\xi(q^{k})\lambda^{k+1/2}$. The standard choice $v_{z}(t_{k})=\frac{z(t_{k+1})-z(t_{k})}{\Delta t}$ ensures that $\xi(\widetilde{q}^{k+1})=z(t_{k+1})+\mathcal{O}(\Delta t^{2})$. However, such a choice is not possible in our context since it would violate the reversibility property (II.8). One simple choice that fulfills the requirements is presented in Section˜II.2.2. Another canonical choice with a meaningful limit in continuous time that satisfies the reversibility condition is
$v_{z}(t_{k})=\frac{z(t_{k+1})-z(t_{k-1})}{2\Delta t},0\leq k\leq K_{T},$
requiring just the auxiliary extension of the position schedule to times $t_{-1}$ and $t_{K_{T}+1}$. To fulfill the requirement that $v_{z}(t_{0})=v_{z}(t_{K_{T}})=0$, this implies that $z(t_{-1})=z(t_{1})$ and $z(t_{K_{T}+1})=z(t_{K_{T}-1})$.

### v3

@@ III.3 Dimer in a solvent
We now apply the algorithm on a model with a non-linear collective variableStraub et al. (1988); Dellago et al. (1999); Lelièvre et al. (2010). It consists of $N$ particles of equal masses $M$ in a two-dimensional box of side length $L$ with periodic boundary conditions. Two of the particles form a dimer; the remaining particles are designated as solvent particles. The solvent particles interact between them and with the dimer particles via the Weeks–Chandler–Andersen (WCA) pair potential, which is a truncated Lennard-Jones potential:
@@
The minima of $V_{\mathrm{D}}$ are at $r=r_{0}$ and $r=r_{0}+2w$, giving rise to a compact state and stretched state with an energy barrier of height $h$. Two configurations for the compact and the stretched state are shown in Fig.˜8.
@@
We ran numerical experiments for parameters $h=w=2$, $r^{*}=\epsilon=1$ and $M=1$ with $N=100$ particles and a box of length $L=15$, in line with previous worksLelièvre et al. (2010, 2012). For the CV sampler we use a Gaussian mixture model with modes of equal weight at $z=0$ and $z=1$ and width $\sigma_{\mathrm{prop}}=0.2$. We reject any unphysical proposals with $z<z_{\mathrm{min}}=-\frac{r_{0}}{2w}$ and $z>z_{\mathrm{max}}=(L/\sqrt{2}-r_{0})/(2w)$ since the distance $|q_{1}-q_{2}|$ is always positive and the periodic boundary conditions (torus coordinates) impose a maximum possible distance.
We compare the performance of the algorithm for different parameter values using a mode jump cost that counts the number of calls to the force required to go from a state with $z<0.1$ to a state with $z>0.9$ or vice versa, in line with the ‘mean residence duration’ used in Ref. Lelièvre et al., 2019.
Once again, we observe a large advantage in using the deterministic algorithm with $\alpha_{1}=0$, whose performance for different values of $\alpha_{2}$ and $\widetilde{v}$ are shown in Fig.˜9, while we see almost no mode switches for the overdamped algorithm within the same computational budget (see Section˜H.4).
In Fig.˜10, we show the free energy profile obtained by running the algorithm for the optimized parameters until convergence together with the free energy in the absence of solvent particles (see Section˜F.3 for its derivation). As one can see, the presence of the solvent particles increases the likelihood of the compact state (lowering its free energy in comparison with the stretched state). The result agrees almost perfectly with a result obtained from thermodynamic integration (TI) (see Ref. Lelièvre et al., 2012 for a theoretical introduction). In thermodynamic integration, the system is sampled under the constraint $\xi(q)=z_{i}$ for a sequential series of CV-values $z_{i}$ on a grid to estimate the mean force. This demonstrates that the inaccuracy of the proposal potential has been corrected by the algorithm.
Note that this model was also considered in the original NCMC paperNilmeier et al. (2011). Contrary to our approach, however, they chose the full dimer configuration as the CV, $\xi(q)=(q_{1},q_{2})^{\top}$, and used a proposal kernel that samples a dimer configuration $\widetilde{Z}_{n+1}$ conditioned on the current configuration $Z_{n}$ with a new distance $|q_{1}-q_{2}|$ but the same overall rotation angle.
@@
Figure 10: Free energy profile for the dimer model obtained from running the CV guided algorithm with the optimal parameters determined in Fig.˜9 with 20 chains and 50,000 iterations, shown alongside results from thermodynamic integration on a grid of 50 values and $10^{6}$ steps each. For comparison, we also show the free energy profile of the dimer without solvent particles, see Section˜F.3, and the energy of the proposal distribution. An energy shift is applied to all of these curves such that they coincide at $z=1$. Error bars are estimated from the standard deviation over parallel MCMC chains of counts of the histogram bins.
@@ III.4 Polymer in a solvent
As a final example, we consider a toy molecular system that consists of a 9-bead polymer embedded in a three-dimensional box with 200 solvent particles, similar to a model introduced in Ref. Tamagnone et al., 2024. The polymer beads and the solvent particles all interact via a Lennard-Jones (LJ) potential. Within the polymer, the nearest and second-nearest neighbor distances (and hence the binding angle) of the polymer beads are softly constrained with a harmonic potential. In addition, a bi-modal potential is imposed on the end-to-end distance of the polymer, giving rise to two stable confirmations in an open and a closed state, see Fig.˜11. Denoting by $\{q_{i}\}_{i=1,\dots,9}$ the polymer coordinates and by $\{q_{i}\}_{i=10,\dots,N}$ the solvent coordinates, the total energy is given by
@@
Figure 12: Simulation of the polymer with a one-dimensional CV (end-to-end distance). Left panel: Rough grid optimization of parameters for the deterministic case $\alpha_{1}=0$, each simulation run with 20 parallel chains and 2,000 iterations. Right panel: Marginals for a run of 20 chains and 100,000 iterations for the optimized parameters $1/\widetilde{v}=180$, $\alpha_{2}=2\times 10^{-4}$ (overall 8% acceptance), shown alongside the CV proposal distribution and reference samples obtained from a biased simulation that has been reweighted.
@@
An obvious choice of CV that resolves the bimodality is the end-to-end distance. Since it is one-dimensional, it is easy to construct a proposal distribution in the form of a Gaussian mixture. For simplicity, we extract the position of the modes and their width from short Langevin simulations on the full system (see Appendix˜G). A rough grid search to optimize the parameters of the algorithm is shown in Fig.˜12, with the best achieved inverse mode jump cost at around $9\times 10^{-5}$. In contrast, the overdamped algorithm with $\alpha_{1}=1$ does not show any transitions within the same parameter range and computational budget, giving an upper limit of the performance of at most $6\times 10^{-7}$; it is thus at least two orders of magnitude worse in performance.
Another possible choice of collective variable in this case is to use the positions of the beads of the entire polymer configuration. Here, a good proposal distribution requires a careful parameterization, for which we use a normalizing flow (see Appendix˜G). Since our primary interest is in the performance of the algorithm, we trained our proposal sampler on already existing reference data. In practice, one could use adaptive schemes in which the CV-sampler is progressively trained while running the MCMC samplerGabrié et al. (2022); Tamagnone et al. (2024). Note that for this choice of CV, the steered dynamics in Step 2 of the algorithm effectively only acts on the solvent particles, since the schedule already imposes the evolution of the entire polymer. Running the algorithm, we again observe that the deterministic dynamics ($\alpha_{1}=0$) gives good results with frequent mode switches while the overdamped version ($\alpha_{1}=1$) does not work at all. Strikingly, for the optimized parameters, we achieve an overall acceptance rate of 99% with this algorithm. Since the proposals are drawn independently, this corresponds almost to the optimum performance with maximum decorrelation within the polymer degrees of freedom. A rough grid search for the deterministic algorithm is shown in Fig.˜13 together with the marginal distribution of the end-to-end distance. Comparing the optimal mode jump cost obtained with the two choices of CV, we see an advantage of around one order of magnitude in using the positions of all polymer beads over the end-to-end distance between first and last bead as a CV. This only accounts for the calls to the force, however, and not for the effort in learning the proposal sampler in CV space.
@@ IV Conclusion and outlook
In this work, we present an MCMC algorithm with non-local proposals in a collective variable space. This CV space may be high-dimensional (and thus classic free energy adaptive biasing methods do not apply), provided a good proposal sampler can still be learned (e.g. by a normalizing flow for CVs of moderately large dimensions). We generalize related algorithms by considering the underdamped Langevin dynamics and a non-linear collective variable and we prove reversibility of the resulting scheme. We provide a detailed description of the algorithm so that it is readily implementable and we illustrate its performance on a number of examples, ranging from a toy low-dimensional example to molecular systems. We demonstrate a significant performance advantage in using Hamiltonian over overdamped steered Langevin dynamics. As mentioned in the introduction, our algorithm enters into frameworks like NCMC and HNMD, but we treat here all technical details arising from a non-linear CV and provide a numerical recipe for the use of general underdamped Langevin dynamics.
There is a tradeoff in this kind of algorithm between the facility of finding a ‘good’ CV and learning a good proposal sampler in that CV space. The system needs to be able to follow the constrained dynamics (Step 2 of the algorithm) and still arrive at a low-energy configuration, for which it is vital that the CV resolves the metastability (see also Appendix B in Ref. Schönle et al., 2025). Evidently, this is easier to achieve allowing the CV to be higher-dimensional. In the (absurd) limiting case where the CV comprises all degrees of freedom of the system, the ‘steering’ is trivially simple. At the same time, the higher-dimensional the CV space becomes, the harder it will be to parameterize and learn an efficient proposal sampler. Our polymer example with the 27-dimensional CV shows how there can be a sweet spot at intermediate dimensionality. Although we do not account for the training effort here, we demonstrate that a normalizing flow can be trained to such high accuracy that it can be used efficiently in the framework of this algorithm. When the proposal sampler has to be learned, one may resort to adaptive frameworks alternating between sampling and training, which have been demonstrated to be effective for normalizing flowsGabrié et al. (2022), assuming that a good initial guess of the mode locations is available.
@@ Appendix A Numerical implementation of the evolution of momenta
For practical purposes, one can reformulate (II.17) (and (II.19) in the same fashion) in the following equivalent form that is more suitable to numerical implementation, see Sec. 5.4.2 in Ref. Lelièvre et al., 2012 for further explanations:
@@ Appendix B Proof of the reversibility of the algorithm
The objective of this section is to prove the reversibility of the algorithm introduced in Section II.2 (namely Theorem˜II.3). This is done in three steps. In Section B.1, we introduce the necessary material, in particular concerning measures on submanifolds associated with the collective variable $\xi$. Then, in Section B.2, we present fundamental properties of the steered dynamics (II.17)-(II.19), and in particular the Jazynski-Crooks equality. This equality is the cornerstone of the proof of reversibility, which is finally provided in Section B.3.
@@ B.1 Introducing notation
This subsection largely uses tools and results from Ref. Lelièvre et al., 2012. Let us introduce a number of quantities which we will need later on. A central quantity of interest is the free energy $F:\mathbb{R}^{\ell}\rightarrow\mathbb{R}$ associated with the target measure $\nu$ and the collective variable $\xi$, defined as (see (II.2) for the definition of $\Sigma(z)$)
$F(z)=-\frac{1}{\beta}\log\int_{\Sigma(z)}\mathcal{Z}_{q}^{-1}e^{-\beta V(q)}\delta_{\xi(q)-z}(\mathrm{d}q).$ (B.1)
@@
The proof relies on Ref. Lelièvre et al., 2012. From Sec. 2.3 therein, we restate the following two identities: By the co-area formula, one can show that
@@
where we denote by $\sigma^{M}_{\Sigma(z)}(\mathrm{d}q)$ the surface measure on $\Sigma(z)$ induced by the scalar product $\langle q,\tilde{q}\rangle_{M}=q^{\top}M\tilde{q}$ on $\mathbb{R}^{d}$, and, for a given $q\in\Sigma(z)$, by $\sigma^{M^{-1}}_{\Sigma_{v_{\xi}(q,\cdot)}(v_{z})}(\mathrm{d}p)$ the surface measure on the affine space $\Sigma_{v_{\xi}(q,\cdot)}(v_{z})$, induced by the scalar product $\langle p,\tilde{p}\rangle_{M^{-1}}=p^{\top}M^{-1}\tilde{p}$ on $\mathbb{R}^{d}$. For more precise definitions of these measures, we refer to Ref. Lelièvre et al., 2010, Sections 3.2.1 and 3.3.2 and the references therein.
@@
Equation (B.5) follows by taking $v_{z}=0$. For the last but one equality in the series of equalities above, we use a computation from Sec. 5.1 in Ref. Lelièvre et al., 2012, namely, for fixed $q$,
@@ B.2 Properties of the steered dynamics
In this section, we state a number of properties of the discretization scheme introduced in (II.17)-(II.19). Since we follow dynamics for the energy modified by the Fixman term, in principle we would need to put a tilde on all associated measures and quantities, but we omit it here for the sake of readability and just consider the dynamics for a general potential $V$.
@@ B.2.1 Forward steered dynamics
since the Gram tensor $G_{M}(q^{k+1})=\nabla\xi(q^{k+1})^{T}M^{-1}\nabla\xi(q^{k+1})$ is invertible by assumption. In contrast, the constraint $(C_{q})$ leads to a non-linear equation for $\lambda^{k+1/2}$, which is solved using Newton’s method. Regarding uniqueness of the solution and convergence of this solver, we require a reversibility property between the forward dynamics and the backward steered dynamics introduced in Section˜B.2.2. This property is explained at the end of Section˜B.2.4, defined by (B.21). Loosely speaking, one needs to have a consistent procedure which gives the same solution when following the forward or backward dynamics.
@@
Unlike the standard RATTLE discretization, our formulation considers time-evolving constraints. Moreover, in contrast to Ref. Lelièvre et al., 2012, Sec. 5.4.1, where such versions of RATTLE are also considered, it uses a separate schedule for the velocity $(v_{z}^{(Z_{n},\widetilde{Z}_{n+1})}(t_{k}))_{0\leq n\leq K_{T}}$, that can in principle be chosen independently of the position schedule $(z^{(Z_{n},\widetilde{Z}_{n+1})}(t_{k}))_{0\leq k\leq K_{T}}$. Still, it is numerically advantageous to use a schedule with a meaningful limit in continuous time, such that (in the limit $\Delta t\to 0$) the position schedule converges to $\xi(q(t))=z(t)$ where $z$ is a fixed continuously-differentiable function and such that the velocity schedule converges to $v_{z}(t)=\dot{z}(t)$. To see this, we write the unconstrained part of the update in (II.18) as
@@
which is followed by a projection with $q^{k+1}=\widetilde{q}^{k+1}+\Delta tM^{-1}\nabla\xi(q^{k})\lambda^{k+1/2}$, where $\lambda^{k+1/2}$ is such that $\xi(q^{k+1})=z(t_{k+1})$. Two examples of velocity schedule that converge to $v_{z}(t)=\dot{z}(t)$ in the continuous-time limit are (1) trivially taking $v_{z}(t_{k})=\dot{z}(t_{k})$ (as in (II.9)-(II.10) in Section˜II.2.2) or (2) the discretization
$v_{z}(t_{k})=\frac{z(t_{k+1})-z(t_{k-1})}{2\Delta t},\,0\leq k\leq K_{T}.$
For this second choice, satisfying the condition (II.6) ($v_{z}(t_{0})=v_{z}(t_{K_{T}})=0$) implies extending the position schedule to times $t_{-1}$ and $t_{K_{T}+1}$ and setting $z(t_{-1})=z(t_{1})$ and $z(t_{K_{T}+1})=z(t_{K_{T}-1})$. For both choices, one has that $\xi(\widetilde{q}^{k+1})=z(t_{k+1})+\mathcal{O}(\Delta t^{2})$ (using that $K_{T}=T/\Delta t=\mathcal{O}(\Delta t^{-1})$, where $T$ is the physical transition time) and thus the requirement $(C_{q})$ is already approximately fulfilled by $\widetilde{q}^{k+1}$, facilitating convergence of the Newton solver for $\lambda^{k+1/2}$. Note, however, that the simplest choice of discretization $v_{z}(t_{k})=\frac{z(t_{k+1})-z(t_{k})}{\Delta t}$ which was used in Ref. Lelièvre et al., 2012, Sec. 5.4.1, is not possible in our context since it would violate the reversibility property (II.8).